OpenStudy (anonymous):

In 1999, the united states produced about 2 million more motor vehicles than Japan. Together they produced about 22 million motor vehicles. How many motor vehicles were produced in each contry

OpenStudy (anonymous):

12 million in the united states, 10 in Japan

OpenStudy (anonymous):

can u show how u got it

OpenStudy (anonymous):

2X=20,000,000 X=20,000,000/2 X=10,000,000 NUMBER OF CARS PRODUCED IN JAPAN 22,000,000-10,000,000=12,000,000 NUMBER OF CARD PRODUCED IN US.

OpenStudy (anonymous):

Call the total number of cars 't', the amount in the US 'u' and the amount in Japan 'j'. We know: \[u=j+2\] and \[j+u=22\]so\[j+j+2=22\]\[2j+2=22\]\[2j=20\]\[j=10\] and \[u=j+2\]\[u=10+2\]\[u=12\] USA produced 12million, Japan produced 10million for a total of 22million.

OpenStudy (anonymous):

thx